# Complexity of equal 0-surgeries

### v2

@@ 4. Mirrors
We compute the signatures of all prime knots with at most $15$ crossings. Whenever the signature is non-vanishing then $K$ and $-K$ are not friends. The signature is vanishing for exactly $80015$ knots. We use SnapPy to check that $353$ of the remaining knots are actually isotopic to their mirrors. Among the remaining $79662$ knots exactly $44$ have non-hyperbolic 00-surgeries. If any of the latter admits an orientation-reversing diffeomorphism, this diffeomorphism is isotopic to one that preserves the JSJ decomposition. In particular, there exist orientation reversing-diffeomorphisms of all JSJ pieces. If a piece of the JSJ decomposition is a Seifert fibered space, then we can use the classification of Seifert fibered spaces that admit orientation-reversing diffeomorphism, see for example Section 8 in [NR78]. Thus we can read-off from the Regina names in Table 2 directly if a Seifert fibered JSJ piece has an orientation-reversing diffeomorphism. If a JSJ decomposition admits a hyperbolic piece then we can use the symmetry group calculations in SnapPy of that piece to check if it admits an orientation-reversing diffeomorphism. By sorting out the 00-surgeries that have a JSJ decomposition with at least one Seifert fibered or hyperbolic piece without orientation-reversing diffeomorphism we are left only with the knot $K13n469$ whose 00-surgery has the JSJ decomposition consisting of the figure eight knot complement $m004$ and the Seifert fibered space $SFS[D:(2,1)(2,-1)]$. Both spaces admit orientation reversing-diffeomorphisms. But since Dunfield’s Regina code does not return the gluing map, we do not directly know if that space admits an orientation-reversing diffeomorphism. But by Remark 4.2, we know that $K13n469$ is actually the first knot in a family of infinitely many knots that are friends with their mirrors.
@@
It remains to handle the chiral knots with vanishing signature and hyperbolic 00-surgeries. For these, we use SnapPy to compute the symmetry groups of the 00-surgeries from which we can read-off if the 00-surgeries admit an orientation-reversing diffeomorphism. It turns out that $K14n3411$ and $K15n64176$ are the only chiral knots that have amphicheiral 00-surgeries.
@@
The same strategy works also for the census knots. Of the $1267$ census knots exactly $139$ have vanishing signatures. Among these $5$ are isotopic to their mirrors. Of the remaining knots, $56$ have non-hyperbolic 00-surgery and $78$ have hyperbolic 00-surgery. Of the non-hyperbolic 00-surgeries we can exclude as above all but two ($v2272$ and $o9\_22951$) to admit an orientation-reversing diffeomorphism. The knot $v2272$ is in fact $K13n469$ and thus is a friend of its mirror by Remark 4.2. The 00-surgery of $o9\_22951$ has JSJ decomposition consisting of the figure eight knot complement $m004$ and the Seifert fibered space $SFS[D:(3,1)(3,-1)]$. In Theorem 9.1 below we show that $o9\_22951$ is a friend of its mirror by constructing an RGB link for that pair.
@@
For the knots with hyperbolic 00-surgeries, we use SnapPy to compute their symmetry groups and deduce that only two admit an orientation-reversing diffeomorphism. The knots $t11462$ and $o9\_22951$ both have hyperbolic 00-surgeries that admit orientation-reversing diffeomorphisms. Thus they are friends with their mirrors.
@@
We also remark that the knots $K14n3411$, $K15n64176$, $t11462$, and $o9\_22951$ are chiral but have amphicheiral 00-surgeries. It follows that these knots necessarily have 00 as a symmetry-exceptional slope [BKM24b], cf. Proposition 8.3, since the symmetry groups of their 00-surgeries are larger than the symmetry groups of the knots. Indeed, we have $\operatorname{Sym}(t11462)=\mathbb{Z}_{2}$ and $\operatorname{Sym}(t11462(0))=D_{4}$. For the other three knots, the symmetry groups are trivial, while their 00-surgeries have symmetry group $\mathbb{Z}_{2}$. ∎
@@ 5. Burton’s and Thistlethwaite’s lists
To prove Theorem 1.1 and for the creation of the census from Theorem 1.4 we need to show that some low crossing knots do not have friends of larger crossing numbers. More precisely, let $K$ be a prime knot with crossing number $c(K)$ at most $9$, we need to show that there exists no prime knot $K^{\prime}$ with crossing number $c(K^{\prime})\leq 25-c(K)$ such that $K$ and $K^{\prime}$ have diffeomorphic 00-surgeries.
For that we first remark that the unknot [KMOS07], the trefoil $K3a1$ [OS19], the figure-eight knot $K4a1$ [OS19], and $K5a1$ [BS22b] do not have any friends. So the first knot, that might have a friend is the $(5,2)$-torus knot $K5a2$. (However, it is conjectured that it has no friend at all.)
@@
For that we run through Burton’s lists of prime knots of at most $19$ crossings [Bur20] and Thistlethwaite’s list of prime knots with $20$ crossings [Thi23], compute their Alexander polynomials and if one of the Alexander polynomials turns out to be the same as the Alexander polynomial of a knot with low-crossing number, we distinguish the 00-surgeries by using the hyperbolicity status, the volume, the Regina names, or the fundamental groups as in Section 2. If these invariants all agree, we use SnapPy and Regina to search for a diffeomorphism between their 00-surgeries. This successfully determines for all possible pairs of knots if they are friends or not.
@@ 6. Census knots
The same strategy as for the low-crossing knots works also for the census knots. But here everything works much faster since there are only $1267$ census knots whose complements can be triangulated by at most $9$ tetrahedra. Moreover, the classification of the exceptional fillings on the census knots was already done by Dunfield [Dun20a], and many of its other invariants were already computed in [ABG+23, BK24, BBK+]. With the same approach, we also classify all census knots that have a friend among the knots with at most $15$ crossings.
@@ 7. Smooth 4-genus
It remains to perform the $4$-genus calculations in Theorem 1.4. For the knots with at most $12$ crossings and for the census knots we can read-off the data from existing databases [LM23, ABG+23, BK24, BBK+]. For the other knots, we search for a ribbon description using the code from [GHMR23]. If this code could not identify a knot as being ribbon, we checked the standard sliceness obstructions, such as the signature, the Fox–Milnor condition, the $\tau$- and $s$-invariant, and the obstruction from [HKL10]. This was enough to determine the sliceness status of all but two knots in Table 1. The two knots with undetermined sliceness status are friends of their mirrors and so they have the same $4$-genus.
To show that all non-slice knots in that table have smooth $4$-genus equal to $1$, we search for a genus-$1$ concordance to a ribbon knot. This is done by searching for a crossing change that yields a knot that is identified by [GHMR23] to be ribbon.
@@ 8. Extending diffeomorphisms over traces and symmetry-exceptional slopes
To show in practice that two friends are $4$-dimensional friends, we describe a 00-surgery diffeomorphism $\varphi\colon K(0)\rightarrow K^{\prime}(0)$ via explicit Kirby moves and then analyze the image of the meridian $\mu$ of $K$ under $\varphi$. This can reveal that $\varphi$ extends to a diffeomorphism of the 00-traces, see for example [Akb77, AJOT13, MP23].
It is often harder to show that two friends are not $4$-dimensional friends. We first recall a result of Boyer [Boy86]. Let $K$ and $K^{\prime}$ be friends and let $\varphi\colon K(0)\rightarrow K^{\prime}(0)$ be a diffeomorphism. We define the parity of $\varphi$ to be the parity of the intersection form of the closed $4$-manifold
@@
Theorem 8.1 (Boyer [Boy86], cf. Theorem 3.7 in [MP23])
@@
We remark that there could exist 00-surgery diffeomorphisms that extend to trace homeomorphisms but not to trace diffeomorphisms. Moreover, there exist examples of friends that have two different 00-surgery homeomorphisms $\varphi,\varphi^{\prime}\colon K(0)\rightarrow K^{\prime}(0)$, such that $\varphi$ is odd and thus does not extend to a homeomorphism of the traces but such that $\varphi^{\prime}$ is even and even extends to a trace diffeomorphism. We refer to Examples 9.5 and 9.6 for concrete examples.
On the other hand, [MP23] describe situations in which every 00-surgery diffeomorphism extends to a trace homeomorphism.
Theorem 8.2 (Manolescu–Piccirillo [MP23])
If the Arf invariant $\operatorname{Arf}(K)$ of $K$ is non-vanishing, then any 00-surgery diffeomorphism from $K(0)$ to any other 00-surgery $K^{\prime}(0)$ extends to a trace homeomorphism.
Note that the Arf invariant of a knot can be computed from its Alexander polynomial and since the Alexander polynomial of a knot is an invariant of its 00-surgery, it follows that this condition is symmetric. We have used SnapPy to compute the Arf invariants of all pairs of friends from our census. The friends with non-vanishing Arf invariants are displayed in Table 5. By Theorem 8.2 these friends have homeomorphic traces.
@@
On the other hand, to show that $K$ and $K^{\prime}$ have non-homeomorphic traces we need to check that all 00-surgery diffeomorphisms $\varphi\colon K(0)\rightarrow K^{\prime}(0)$ are odd. In the following, we present a criterium to show that any 00-surgery diffeomorphism of two friends is odd and thus the friends are not $4$-dimensional friends.
First, we recall that if $K$ is a hyperbolic knot, then Thurston’s hyperbolic Dehn filling theorem tells us that for all but finitely many slopes $r$, the Dehn filling $K(r)$ of $K$ with slope $r$ is again hyperbolic [Thu79]. (The finitely many slopes for which $K(r)$ is not hyperbolic are called exceptional.) Moreover, if the slope $r$ is sufficiently long then the core of the newly glued-in solid torus is the shortest geodesic of $K(r)$. By Mostow rigidity, any diffeomorphism of a hyperbolic manifold is isotopic to a unique isometry. Since an isometry preserves the shortest geodesic, it follows that for sufficiently large $r$, any symmetry of $K(r)$ restricts to a symmetry of $K$. If there exists a symmetry of $K(r)$ that does not restrict to a symmetry of $K$ then $r$ is called symmetry-exceptional slope. In particular, it follows that if $K(0)$ has trivial symmetry group, then 00 is not a symmetry-exceptional slope. We refer to [BKM24b] for further discussion. We also mention that there exist computational methods to check if a given slope is a symmetry-exceptional slope [BKM24b].
@@
Let $K$ and $K^{\prime}$ be friends. If $K$ is hyperbolic and 00 is not a symmetry-exceptional slope of $K$, then all 00-surgery diffeomorphisms $K(0)\rightarrow K^{\prime}(0)$ have the same parity.
@@
Opposed to Theorem 8.2, the condition in Proposition 8.3 is not symmetric. There exist friends $K$ and $K^{\prime}$ such that 00 is a symmetry-exceptional slope for $K$ but not for $K^{\prime}$, see for example Table 6.
@@
Let $\varphi,\varphi^{\prime}\colon K(0)\rightarrow K^{\prime}(0)$ be two 00-surgery diffeomorphisms. We write $f\colon K(0)\rightarrow K(0)$ for the unique diffeomorphism such that $\varphi^{\prime}=\varphi\circ f$. By Mostow rigidity, $f$ is isotopic to a unique isometry and thus we can see $f$ as an element in the symmetry group of $K(0)$. Since we assume that 00 is not a symmetry-exceptional slope of $K$, we conclude that $f$ restricts to an isometry of the complement $S^{3}\setminus K$ of $K$. By the solution of the knot complement problem [GL89], such a symmetry sends a 00-framed meridian $\mu$ to itself. We conclude that $\varphi(\mu_{K})$ is isotopic as framed knot to $\varphi^{\prime}(\mu_{K})$. Now we observe, that we obtain the closed $4$-manifold $Z_{\varphi}=X(K^{\prime})\cup_{\varphi}-X(K)$ by attaching $2$-handles to $D^{4}$ along the 00-framed $K^{\prime}$ and the image of the 00-framed meridian $\mu_{K}$ of $K$ under $\varphi$ and capping of with a $4$-handle. But since $\varphi(\mu_{K})$ is isotopic to $\varphi^{\prime}(\mu_{K})$, the closed $4$-manifolds $Z_{\varphi}$ and $Z_{\varphi^{\prime}}$ are diffeomorphic and thus $\varphi$ and $\varphi^{\prime}$ have the same parity. ∎
For applying the obstruction from Proposition 8.3 we classify the symmetry-exceptional 00-surgeries in our census.
@@
The classification of exceptional and symmetry-exceptional 00-surgeries and their symmetry groups of the friends from Table 1 is as displayed in Table 6.
@@
We use the functions in SnapPy to compute the symmetry groups of the knots in our census and to verify the hyperbolicity of their 00-surgeries. If the latter fails, we use Dunfield’s Regina recognition code [Dun20a, FPS24] to recognize the 00-surgery as a non-hyperbolic manifold confirming that it is exceptional. For the hyperbolic 00-surgeries we use the strategies explained in [BKM24b] to compute their symmetry groups. This process is not always guaranteed to terminate, but for the knots in the census, it terminates quickly and yields verified results. ∎
Table 6. This table displays the symmetry groups of the knots from Table 1 and the types and symmetry groups of the 00-surgeries.
@@
$K6a1$ $D_{4}$ $7819nh\_$ 78 00 hyperbolic $D_{4}$
$K9n4$ $\mathbb{Z}_{2}$ $o9\_34801$ 00 hyperbolic $\mathbb{Z}_{2}^{2}$
$K10n10$ 00 $-t11532$ 00 hyperbolic 00
$K10n10$ 00 $o9\_43446$ 00 hyperbolic 00
$K10n13$ $\mathbb{Z}_{2}$ $o9\_34818$ 00 $SFS\,[S^{2}:(2,1)(5,1)(10,-7)]$ at least $\mathbb{Z}_{2}$
@@
$K11n49$ $\mathbb{Z}_{2}$ $o9\_42735$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K11n116$ 00 $-t12607$ 00 hyperbolic $\mathbb{Z}_{2}$
$K12n121$ $\mathbb{Z}_{2}$ $-t11441$ $\mathbb{Z}_{2}$ $SFS\,[D^{2}:(2,1)(2,1)]\cup_{m}$ at least $\mathbb{Z}_{2}$
@@
$K12n309$ $\mathbb{Z}_{2}$ $-K14n14254$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}$
$K12n318$ 00 $-o9\_39433$ 00 hyperbolic 00
$K13n469$ $\mathbb{Z}_{2}$ $-K13n469$ $\mathbb{Z}_{2}$ $JSJ\big{(}SFS\,[D^{2}:(2,1)(2,-1)]\cup m004\big{)}$ at least $\mathbb{Z}_{2}$
$K13n572$ $\mathbb{Z}_{2}$ $-K15n89587$ 00 hyperbolic $\mathbb{Z}_{2}^{2}$
$K13n1021$ $\mathbb{Z}_{2}$ $K15n101402$ 00 hyperbolic $\mathbb{Z}_{2}^{2}$
$K13n1021$ $\mathbb{Z}_{2}$ $-o9\_34908$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K13n2527$ 00 $-K15n9379$ $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K13n3158$ 00 $o9\_41909$ 00 hyperbolic $\mathbb{Z}_{2}$
$K14n3155$ $\mathbb{Z}_{2}$ $-K14n3155$ $\mathbb{Z}_{2}$ hyperbolic $D_{4}$

### v3

@@ 4. Mirrors
We compute the signatures of all prime knots with at most $15$ crossings. Whenever the signature is non-vanishing then $K$ and K-K are not friends. The signature is vanishing for exactly $80015$ knots. We use SnapPy to check that $353$ of the remaining knots are actually isotopic to their mirrors. Among the remaining $79662$ knots exactly $44$ have non-hyperbolic 0-surgeries. If any of the latter admits an orientation-reversing diffeomorphism, this diffeomorphism is isotopic to one that preserves the JSJ decomposition. In particular, there exist orientation-reversing diffeomorphisms of all JSJ pieces. If a piece of the JSJ decomposition is a Seifert fibered space, then we can use the classification of Seifert fibered spaces that admit orientation-reversing diffeomorphisms, see for example Section 8 in [41]. Thus we can read off from the Regina names in Table 2 directly if a Seifert fibered JSJ piece has an orientation-reversing diffeomorphism. If a JSJ decomposition admits a hyperbolic piece then we can use the symmetry group calculations in SnapPy of that piece to check if it admits an orientation-reversing diffeomorphism. By sorting out the 0-surgeries that have a JSJ decomposition with at least one Seifert fibered or hyperbolic piece without orientation-reversing diffeomorphism, we are left only with the knot $K13n469$ whose 0-surgery has the JSJ decomposition consisting of the figure eight knot complement $m004$ and the Seifert fibered space SFS[D:$(2,1)$]SFS[D:(2,1)(2,-1)]. Both spaces admit orientation-reversing diffeomorphisms. But since Dunfield’s Regina code does not return the gluing map, we do not directly know if that space admits an orientation-reversing diffeomorphism. But by Remark 4.2, we know that $K13n469$ is actually the first knot in a family of infinitely many knots that are friends with their mirrors.
@@
It remains to handle the chiral knots with vanishing signature and hyperbolic 0-surgeries. For these, we use SnapPy to compute the symmetry groups of the 0-surgeries from which we can read off if the 0-surgeries admit an orientation-reversing diffeomorphism. It turns out that $K14n3411$ and $K15n64176$ are the only chiral knots that have amphicheiral 0-surgeries.
@@
The same strategy works also for the census knots. Of the $1267$ census knots exactly $139$ have vanishing signatures. Among these, $5$ are isotopic to their mirrors. Of the remaining knots, $56$ have non-hyperbolic 0-surgery and $78$ have hyperbolic 0-surgery. Of the non-hyperbolic 0-surgeries we can exclude as above all but two ($v2272$ and o9_22951o9\_22951) to admit an orientation-reversing diffeomorphism. The knot $v2272$ is in fact $K13n469$ and thus is a friend of its mirror by Remark 4.2. The 0-surgery of o9_22951o9\_22951 has JSJ decomposition consisting of the figure eight knot complement $m004$ and the Seifert fibered space SFS[D:$(3,1)$]SFS[D:(3,1)(3,-1)]. In Theorem 9.1 below we show that o9_22951o9\_22951 is a friend of its mirror by constructing an RBG link for that pair.
@@
For the knots with hyperbolic 0-surgeries, we use SnapPy to compute their symmetry groups and deduce that only two admit an orientation-reversing diffeomorphism. The knots $t11462$ and o9_22951o9\_22951 both have hyperbolic 0-surgeries that admit orientation-reversing diffeomorphisms. Thus they are friends with their mirrors.
@@
We also remark that the knots $K14n3411$, $K15n64176$, $t11462$, and o9_22951o9\_22951 are chiral but have amphicheiral 0-surgeries. It follows that these knots necessarily have 0 as a symmetry-exceptional slope [9], cf. Proposition 8.3, since the symmetry groups of their 0-surgeries are larger than the symmetry groups of the knots. Indeed, we have $\operatorname{Sym}(t11462)=\mathbb{Z}_{2}$ and $\operatorname{Sym}(t11462(0))=D_{4}$. For the other three knots, the symmetry groups are trivial, while their 0-surgeries have symmetry group $\mathbb{Z}_{2}$. ∎
@@ 5. Burton’s and Thistlethwaite’s lists
To prove Theorem 1.1 and for the creation of the census from Theorem 1.4 we need to show that some low crossing knots do not have friends of larger crossing numbers. More precisely, let $K$ be a prime knot with crossing number $c(K)$ at most $9$, we need to show that there exists no prime knot $K^{\prime}$ with crossing number $c(K^{\prime})\leq 25-c(K)$ such that $K$ and $K^{\prime}$ have diffeomorphic 0-surgeries.
For that we first remark that the unknot [35], the trefoil $K3a1$ [45], the figure-eight knot $K4a1$ [45], and $K5a1$ [13] do not have any friends. So the first knot, that might have a friend is the $(5,2)$-torus knot $K5a2$. (However, it is conjectured that it has no friend at all.)
@@
For that we run through Burton’s lists of prime knots of at most $19$ crossings [17] and Thistlethwaite’s list of prime knots with $20$ crossings [52], compute their Alexander polynomials and if one of the Alexander polynomials turns out to be the same as the Alexander polynomial of a knot with low-crossing number, we distinguish the 0-surgeries by using the hyperbolicity status, the volume, the Regina names, or the fundamental groups as in Section 2. If these invariants all agree, we use SnapPy and Regina to search for a diffeomorphism between their 0-surgeries. This successfully determines for all possible pairs of knots if they are friends or not.
@@ 6. Census knots
The same strategy as for the low-crossing knots works also for the census knots. But here everything works much faster since there are only $1267$ census knots whose complements can be triangulated by at most $9$ tetrahedra. Moreover, the classification of the exceptional fillings on the census knots was already done by Dunfield [22], and many of its other invariants were already computed in [7, 12, 8]. With the same approach, we also classify all census knots that have a friend among the knots with at most $15$ crossings.
@@ 7. Smooth 4-genus
It remains to perform the $4$-genus calculations in Theorem 1.4. For the knots with at most $12$ crossings and for the census knots we can read off the data from existing databases [38, 7, 12, 8]. For the other knots, we search for a ribbon description using the code from [30]. If this code could not identify a knot as being ribbon, we checked the standard sliceness obstructions, such as the signature, the Fox–Milnor condition, the $\tau$- and $s$-invariant, and the obstruction from [32]. This was enough to determine the sliceness status of all but the two knots (o9_22951o9\_22951 and o9_41058o9\_41058) in Table 1. The two knots with undetermined sliceness status are friends of their mirrors and so they have the same $4$-genus.
To show that all non-slice knots and the two knots with undetermined sliceness status in that table have smooth $4$-genus at most $1$, we search for a genus-$1$ concordance to a ribbon knot. This is done by searching for a crossing change that yields a knot that is identified by [30] to be ribbon.
@@ 8. Extending diffeomorphisms over traces and symmetry-exceptional slopes
To show in practice that two friends are $4$-dimensional friends, we describe a 0-surgery diffeomorphism $\varphi\colon K(0)\rightarrow K^{\prime}(0)$ via explicit Kirby moves and then analyze the image of the meridian $\mu_{K}$ of $K$ under $\varphi$. This can reveal that $\varphi$ extends to a diffeomorphism of the 0-traces, see for example [6, 2, 39].
It is often harder to show that two friends are not $4$-dimensional friends. We first recall a result of Boyer [15]. Let $K$ and $K^{\prime}$ be friends and let $\varphi\colon K(0)\rightarrow K^{\prime}(0)$ be a diffeomorphism. We define the parity of $\varphi$ to be the parity of the intersection form of the closed $4$-manifold
@@
Theorem 8.1 (Boyer [15], cf. Theorem 3.7 in [39])
@@
We remark that there could exist 0-surgery diffeomorphisms that extend to trace homeomorphisms but not to trace diffeomorphisms. Moreover, there exist examples of friends that have two different 0-surgery homeomorphisms $\varphi,\varphi^{\prime}\colon K(0)\rightarrow K^{\prime}(0)$, such that $\varphi$ is odd and thus does not extend to a homeomorphism of the traces but such that $\varphi^{\prime}$ is even and even extends to a trace diffeomorphism. We refer to Examples 9.5 and 9.6 for concrete examples.
On the other hand, [39] describe situations in which every 0-surgery diffeomorphism extends to a trace homeomorphism.
Theorem 8.2 (Manolescu–Piccirillo [39])
If the Arf invariant $\operatorname{Arf}(K)$ of $K$ is non-vanishing, then any 0-surgery diffeomorphism from $K(0)$ to any other 0-surgery $K^{\prime}(0)$ extends to a trace homeomorphism.
Note that the Arf invariant of a knot can be computed from its Alexander polynomial and since the Alexander polynomial of a knot is an invariant of its 0-surgery, it follows that this condition is symmetric. We have used SnapPy to compute the Arf invariants of all pairs of friends from our census. The friends with non-vanishing Arf invariants are displayed in Table 5. By Theorem 8.2 these friends have homeomorphic traces.
@@
On the other hand, to show that $K$ and $K^{\prime}$ have non-homeomorphic traces we need to check that all 0-surgery diffeomorphisms $\varphi\colon K(0)\rightarrow K^{\prime}(0)$ are odd. In the following, we present a criterium to show that any 0-surgery diffeomorphism of two friends is odd and thus the friends are not $4$-dimensional friends.
First, we recall that if $K$ is a hyperbolic knot, then Thurston’s hyperbolic Dehn filling theorem tells us that for all but finitely many slopes $r$, the Dehn filling $K(r)$ of $K$ with slope $r$ is again hyperbolic [53]. (The finitely many slopes for which $K(r)$ is not hyperbolic are called exceptional.) Moreover, if the slope $r$ is sufficiently long then the core of the newly glued-in solid torus is the shortest geodesic of $K(r)$. By Mostow rigidity, any diffeomorphism of a hyperbolic manifold is isotopic to a unique isometry. Since an isometry preserves the shortest geodesic, it follows that for sufficiently large $r$, any symmetry of $K(r)$ restricts to a symmetry of $K$. If there exists a symmetry of $K(r)$ that does not restrict to a symmetry of $K$ then $r$ is called symmetry-exceptional slope. In particular, it follows that if $K(0)$ has trivial symmetry group, then 0 is not a symmetry-exceptional slope. We refer to [9] for further discussion. We also mention that there exist computational methods to check if a given slope is a symmetry-exceptional slope [9].
@@
Let $K$ and $K^{\prime}$ be friends. If $K$ is hyperbolic and 0 is not a symmetry-exceptional slope of $K$, then all 0-surgery diffeomorphisms $K(0)\rightarrow K^{\prime}(0)$ have the same parity.
@@
Opposed to Theorem 8.2, the condition in Proposition 8.3 is not symmetric. There exist friends $K$ and $K^{\prime}$ such that 0 is a symmetry-exceptional slope for $K$ but not for $K^{\prime}$, see for example Table 6.
@@
Let $\varphi,\varphi^{\prime}\colon K(0)\rightarrow K^{\prime}(0)$ be two 0-surgery diffeomorphisms. We write $f\colon K(0)\rightarrow K(0)$ for the unique diffeomorphism such that $\varphi^{\prime}=\varphi\circ f$. By Mostow rigidity, $f$ is isotopic to a unique isometry and thus we can see $f$ as an element in the symmetry group of $K(0)$. Since we assume that 0 is not a symmetry-exceptional slope of $K$, we conclude that $f$ restricts to an isometry of the complement $S^{3}\setminus K$ of $K$. By the solution of the knot complement problem [29], such a symmetry sends a 0-framed meridian $\mu_{K}$ of $K$ to itself. We conclude that $\varphi(\mu_{K})$ is isotopic as framed knot to $\varphi^{\prime}(\mu_{K})$. Now we observe, that we obtain the closed $4$-manifold $Z_{\varphi}=X(K^{\prime})\cup_{\varphi}-X(K)$ by attaching $2$-handles to $D^{4}$ along the 0-framed $K^{\prime}$ and the image of the 0-framed meridian $\mu_{K}$ of $K$ under $\varphi$ and capping of with a $4$-handle. But since $\varphi(\mu_{K})$ is isotopic to $\varphi^{\prime}(\mu_{K})$, the closed $4$-manifolds $Z_{\varphi}$ and $Z_{\varphi^{\prime}}$ are diffeomorphic and thus $\varphi$ and $\varphi^{\prime}$ have the same parity by Theorem 8.1. ∎
For applying the obstruction from Proposition 8.3 we classify the symmetry-exceptional 0-surgeries in our census.
@@
The classification of exceptional and symmetry-exceptional 0-surgeries and their symmetry groups of the friends from Table 1 is as displayed in Table 6.
@@
We use the functions in SnapPy to compute the symmetry groups of the knots in our census and to verify the hyperbolicity of their 0-surgeries. If the latter fails, we use Dunfield’s Regina recognition code [22, 26] to recognize the 0-surgery as a non-hyperbolic manifold confirming that it is exceptional. For the hyperbolic 0-surgeries we use the strategies explained in [9] to compute their symmetry groups. This process is not always guaranteed to terminate, but for the knots in the census, it terminates quickly and yields verified results. ∎
Table 6. This table displays the symmetry groups of the knots from Table 1 and the types and symmetry groups of the 0-surgeries.
@@
$K6a1$ $D_{4}$ 19nh_7819nh\_78 0 hyperbolic $D_{4}$
$K9n4$ $\mathbb{Z}_{2}$ o9_34801o9\_34801 0 hyperbolic $\mathbb{Z}_{2}^{2}$
$K10n10$ 0 t11532-t11532 0 hyperbolic 0
$K10n10$ 0 o9_43446o9\_43446 0 hyperbolic 0
$K10n13$ $\mathbb{Z}_{2}$ o9_34818o9\_34818 0 $SFS\,[S^{2}:(2,1)(5,1)(10,-7)]$ at least $\mathbb{Z}_{2}$
@@
$K11n49$ $\mathbb{Z}_{2}$ o9_42735o9\_42735 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K11n116$ 0 t12607-t12607 0 hyperbolic $\mathbb{Z}_{2}$
$K12n121$ $\mathbb{Z}_{2}$ t11441-t11441 $\mathbb{Z}_{2}$ $SFS\,[D^{2}:(2,1)(2,1)]\cup_{m}$ at least $\mathbb{Z}_{2}$
@@
$K12n309$ $\mathbb{Z}_{2}$ K14n14254-K14n14254 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}$
$K12n318$ 0 o9_39433-o9\_39433 0 hyperbolic 0
$K13n469$ $\mathbb{Z}_{2}$ K13n469-K13n469 $\mathbb{Z}_{2}$ $JSJ\big(SFS\,[D^{2}:(2,1)(2,-1)]\cup m004\big)$ at least $\mathbb{Z}_{2}$
$K13n572$ $\mathbb{Z}_{2}$ K15n89587-K15n89587 0 hyperbolic $\mathbb{Z}_{2}^{2}$
$K13n1021$ $\mathbb{Z}_{2}$ $K15n101402$ 0 hyperbolic $\mathbb{Z}_{2}^{2}$
$K13n1021$ $\mathbb{Z}_{2}$ o9_34908-o9\_34908 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K13n2527$ 0 K15n9379-K15n9379 $\mathbb{Z}_{2}$ hyperbolic $\mathbb{Z}_{2}^{2}$
$K13n3158$ 0 o9_41909o9\_41909 0 hyperbolic $\mathbb{Z}_{2}$
$K14n3155$ $\mathbb{Z}_{2}$ K14n3155-K14n3155 $\mathbb{Z}_{2}$ hyperbolic $D_{4}$